Diffusion Morning Report 05/18/09

Ted Berg tberg at stanford.edu
Mon May 18 07:12:14 PDT 2009

Good Morning All,
    Looking OK this morning. Tylan1 had MFS alarms and no oxide-found 
Manual valve off. Tylan 3 had alarms during TLC clean.TEOS 2 has a leak 
again. Tylan LTO some temps are a bit off.  B2H6 for diffusion is back 
and will be connected soon.Thermco LTO testing.  No other issues at this 

  	Date: 	5/18/09 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	yellow 	MFS alarms 	  	  	 
Tylan 2 	up 	poor uniformity 	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	up 	  	  	  	 
Tylan 6-POCl3 	up 	  	  	  	 
Tylan 7-TEOS2 	down 	leak 	3.8/30 	  	 
Tylan 8-SiGe 	up 	  	15.82/30 	  	 
Tylan 9-Poly 	up 	  	21.34/30 	  	 
Tylan 10-Nitride 	up 	  	15.54/30 	  	 
Tylan 11-LTO 	yellow 	some temps off abit 	5.2/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	up 	  	  	  	 
Thermco Nitride1 	up 	  	  	  	 
Thermco bpsg 	future 	testing for qual 	  	  	 
TEL Spa 	yellow 	  	  	  	 
Burn Box 	up 	  	   	  	 
Ebeam 	up 	  	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	  	A on line 525 psi 	 changeover working well  	  	 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20090518/e7f4473c/attachment.html>

More information about the diffusion mailing list